USE ONE SAMPLE Z TEST 


A researcher reports that the average salary of College Deans is more than P63,000. A sample , test the claim that the College Deans earn more than P63,000 a month. The standard deviation of the population is P5,250.


Given:


STEP 1: STATE THE HYPOTHESIS AND IDENTIFY THE CLAIM.


STEP 2: THE LEVEL OF SIGNIFICANCE 


STEP 3: THE Z CRITICAL VALUE


STEP 4: COMPUTE THE ONE SAMPLE Z TEST VALUE 


STEP 5: DECISION RULE


STEP 6: CONCLUSION


1
Expert's answer
2021-05-31T19:29:44-0400

Given:

"\mu=P63,000\\n=35\\\bar x=P65,700\\\sigma=P5,250"


STEP 1: State the hypothesis and identify them

"H_0:\mu \leq P63,000\\H_1:\mu>P63,000," claim(One-tailed test)


STEP2: The level of significance is "\alpha=0.05"


STEP3: Determine the critical value using the table

"z_{critical}=+1.645"


STEP4: Compute the one sample z test value using the formula "z=\dfrac{\bar x -\mu}{\sigma/\sqrt n}"


"z_{computed}=\dfrac{\bar x -\mu}{\sigma/\sqrt n}=\dfrac{65,700-63,000}{5250/\sqrt {35}}=(2,700)[\dfrac{\sqrt{35}}{5250}]\\\ \\z_{computed}=3.043"


STEP5: Decision rule. Compare the computed and critical value of z

"\because |3.043|>|1.645|\implies z_{computed}>z_{critical}"


So, we "reject\ H_0\ and \ accept\ H_1"



STEP6: Conclusion:

There is evidence to support the claim that the monthly salary of the college dean is more than P63,000